Today at CRC – 7th March 2016

Bob Stark reports:-


A short report and no pictures. The spring like weather let the gardeners Mike, Dick, Ian and Lynne get out to tidy up the garden beds and pine needles on the ramp and Platform 1. The Ledamun brother have given our grassed areas adjacent to the main station entrance a light haircut whilst managing to avoid the daffodils.
Bob had been “shopping” this morning around various hardware/electrical outlets so that Mick would be able to move the lighting project forward. We now have most of bolt on goodies to put inside the lighting control cabinet on Pl 2 and Mick fitted these. We now only lack a 2nd consumer unti for the 13A power supply into the cabinet and that is now on order. Mick also fitted the key locks that arrived for the cabinet to keep prying fingers away from the electricity and carried on with the detective work to find out how and where the old Pl 2 running in Board lighting circuit is activated. Ros ran our carpet shampoo  machine over the Booking Hall carpet and the colour of the water extracted resembled that of the Nile in flood. Roger continued his remedial attack on the exterior of the Lamp Hut, descaling the old paint and priming with new. Dave Griffin continued his resolution of the few behind the scenes issues that arose over the first weekend of operations and we are hopeful that the snags have now been seen to.
If I have missed out anyone – apologies.

Leave a Reply

Please log in using one of these methods to post your comment: Logo

You are commenting using your account. Log Out /  Change )

Google photo

You are commenting using your Google account. Log Out /  Change )

Twitter picture

You are commenting using your Twitter account. Log Out /  Change )

Facebook photo

You are commenting using your Facebook account. Log Out /  Change )

Connecting to %s